OpenStudy (anonymous):

Multiply. ^3sqrt2 * ^3sqrt4

OpenStudy (anonymous):

2

OpenStudy (anonymous):

\[\sqrt[3]{2}\times \sqrt[3]{4}=\sqrt[3]{8}=2\]

OpenStudy (anonymous):

just ot make sure i got your question right is it: \[3\sqrt{2} \times 3\sqrt{4}\] ?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

the way satellite wrote the problem was right